OpenStudy (anonymous):

Evaluate the following expression using the values given: Find 3x2 - y3 - y3 - z if x = 3, y = -2, and z = -5.

OpenStudy (whpalmer4):

\[3x^2-y^3-y^3-z\]Is that the correct expression? I'm suspicious of the repeated \(y^3\) term in there...

OpenStudy (kamibug):

In the case that the equation is as @whpalmer4 sates above, the steps for solving would be as follows. Substitute the given values into the equation. 3(3)^2 - (-2)^3 - (-2)^3 - (-5) 9^2 - 4 - 4 +5 81 -8 +5 81 - 3 78

OpenStudy (whpalmer4):

Mmm…not quite… \[3x^2-y^3-y^3-z\]\[3(3)^2-y^3-y^3-z=27-y^3-y^3-z\]\[27-(-2)^3-(-2)^3-z = 27-(-8)-(-8)-z = 27+8+8-z = 43-z \]\[43-z = 43-(-5) = 48\]

OpenStudy (whpalmer4):

I did the substitutions one by one so you could see them a bit more clearly.

OpenStudy (kamibug):

My mistake was doing y squared instead of cubed.